OpenStudy (anonymous):

A sample of nitrogen (Mr = 28.0) has a mass of 10 g and a volume of 8 L. Which of the following values can be calculated for the nitrogen sample from these data? I moles II % w/v III density A I only B I and II only C I and III only D I, II and III hello, i was wondering what the answer to this question should be cause in the worked solution (funny enough didnt give any explaination of how they got their ans) got B but the back of the book said C and i said D??? thank you

OpenStudy (somy):

mole concentration density

OpenStudy (somy):

mole=mass/Mr --------------- u both mass and Mr concentration= mole/volume --------- u have u can find mole and thus concentration density=mass/volume ------------you have both mass and volume

OpenStudy (somy):

btw what's % w/v ?

OpenStudy (somy):

your book is right tho it should be C

OpenStudy (somy):

http://dl.clackamas.edu/ch105-04/wtvolpct.htm according to this i see that for %w/v to find you need, solute and solvent, so that u can have Solution BUT u have only Nitrogen

OpenStudy (anonymous):

thank you for helping, very much appreciated :) but can't the mass of nitrogen be the mass of solute?

OpenStudy (somy):

but you do not have solvent right?

OpenStudy (somy):

the volume that you are supposed to use in that formula is Volume of Solution if there is no Solvent, then how can u has Solution?

OpenStudy (anonymous):

isnt the solvent 8L?

OpenStudy (somy):

Noooo lol 8 L is volume of Nitrogen itself :)

OpenStudy (somy):

it could be volume of Solution if this volume of N was with for example mixed with water (water acts as solvent)

OpenStudy (somy):

but its volume of N alone

OpenStudy (somy):

have you learnt about dilution?

OpenStudy (anonymous):

yes i have c1v1 = c2v2

OpenStudy (somy):

see in dilution, its solute (which can be anything) and then u just add water to it to make up some amount of volume in that case this volume that u made up is volume of Solution

OpenStudy (anonymous):

thank you :) i have another question though can u please help me here as well? More energy is required to vibrate: A C–C compared to C=C B C–Cl compared to C–Br C C–O compared to C–H D C=C compared to C≡C shouldn't the answer be B, the back of book says C?

OpenStudy (somy):

well A and D are just too wrong, u understand why right?

OpenStudy (anonymous):

yes cause triple bond is stronger than double bond which is stronger than single bonds

OpenStudy (somy):

good :) are you familiar with Electronegativity concept?

OpenStudy (anonymous):

yes F most electronegative, electron attracting ability

OpenStudy (somy):

okay so Cl is more electronegative then Br right?

OpenStudy (anonymous):

yes

OpenStudy (somy):

but Br is just below Cl right

OpenStudy (somy):

Cl electronegativity values is 3.0 and Br is 2.8 and thus the difference is only 0.2 right?

OpenStudy (somy):

electronegativity value for H is 2.1 and for O is 3.5 difference is 1.4

OpenStudy (somy):

so what do you think which one requires highest amount of energy?

OpenStudy (anonymous):

isnt it between c and cl or c and br not between cl and br?

OpenStudy (somy):

C are going to stay same

OpenStudy (somy):

so u focus on the one that changes

OpenStudy (somy):

alright if u don't get it this way lets go with the way you get it

OpenStudy (anonymous):

sorry

OpenStudy (somy):

C is in group 4

OpenStudy (somy):

u do know that electronegativity increases across the period right?

OpenStudy (anonymous):

yes

OpenStudy (somy):

so anything before C is less electronegative then C and anything after C is more electronegative then C

OpenStudy (somy):

do u agree to this?

OpenStudy (anonymous):

yes

OpenStudy (somy):

next, u do know that electronegativity increases as u go UP the group from the period 7 to period 1 going up it increases

OpenStudy (anonymous):

yes

OpenStudy (somy):

okay cool now look the bond is C-Cl and C-Br Cl is is below and to the right of C and Br is more below but same amount to the right

OpenStudy (somy):

so that means it harder to break C-Cl bond the C-Br bond right?

OpenStudy (anonymous):

yes so the ans is b?

OpenStudy (somy):

now lets go back to electronegativities C - 2.5 Cl- 3.0 C-2.5 Br- 2.8

OpenStudy (somy):

u can understand why Cl is stronger right?

OpenStudy (somy):

no, it is C

OpenStudy (somy):

C- 2.5 O- 3.5 C- 2.5 H- 2.1

OpenStudy (anonymous):

cause i always thought that we should look at it with the mass, the lighter the mass the harder to break the bond

OpenStudy (somy):

difference in electronegativities between C-Cl is 3.0-2.5= 0.5 for C- Br 2.8-2.5= 0.3 and u see that difference in C-Cl is bigger thus its stronger

OpenStudy (somy):

no it's actually all about electronegativity

OpenStudy (somy):

now when u look at C- O the difference is 3.5-2.5= 1 and for C-H its 2.1-2.5= -0.4

OpenStudy (somy):

now tell me difference between 0.5 and 0.3 OR difference between 1 and -0.4 is greater?

OpenStudy (anonymous):

1

OpenStudy (somy):

as u see the difference between C-H and C-O bond is way greater then in C-Cl and C-Br

OpenStudy (somy):

thats why more energy is needed to break C-O then C-Cl

OpenStudy (somy):

i hope i didn't confuse u even more, if i did please tell me

OpenStudy (anonymous):

nope i got it now thank you i know it takes more to break c-o than c - cl but it takes more energy to break c-h than c-o right?

OpenStudy (somy):

yes thats right! u compare the differences between elcetronegativities

OpenStudy (somy):

difference between C- O and C-H was greater the difference between C-Cl and C-Br

OpenStudy (somy):

thus C would be your right answer

OpenStudy (anonymous):

yes i know... but just a side question: c-h requires more energy to break than c-o?

OpenStudy (somy):

no no no lol C-O requires more energy

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (somy):

you are welcome :)